what is the quotient of 22 divided by 8 and put the mixed number in simplest form.






Thank you

Answer by stanbon(75887)

what is the quotient of 22 divided by 8
and put the mixed number in simplest form.
----------------
22/8 = 16/8 + 6/8
= 2 + 3/4
= 2 3/4
======================================
Cheers,
Stan H.